OpenStudy (anonymous):

tan ^{2}theta+(1-sqrt{3})tan theta=sqrt{3} [0º≤θ<360º] Find Θ.

OpenStudy (anonymous):

\[\tan^{2}\theta+(1-\sqrt{3})\tan \theta=\sqrt{3}\]

OpenStudy (agent0smith):

Notice it looks just a little like a quadratic equation... subtract sqrt3 from both sides \[\large \tan^{2}\theta+(1-\sqrt{3})\tan \theta - \sqrt{3} = 0\] \[\large (\tan\theta)^2+(1-\sqrt{3})\tan \theta - \sqrt{3} = 0\] Now to make it more obvious, let \[x=\tan \theta\] \[\large x^2 + (1-\sqrt{3})x - \sqrt3 = 0\] Keep in mind that 1-sqrt3 is really just a constant, so we now have ax^2+bx+c=0.

OpenStudy (anonymous):

does it have any other method except this?

OpenStudy (anonymous):

how about common factors?

OpenStudy (agent0smith):

Well, you can factor it now that it's a quadratic.

OpenStudy (anonymous):

\[x=-1/\sqrt{3}\]

OpenStudy (anonymous):

θ=60/-45(rej.)

OpenStudy (agent0smith):

\[\large x^2 + (1-\sqrt{3})x - \sqrt3 = 0\] you can do it this way by distributing: \[\large x^2 + x-\sqrt{3} x - \sqrt3 = 0 \] then factor: \[\large x(x + 1)-\sqrt{3}(x - 1) = 0 \] \[\large (x-\sqrt{3})(x + 1) = 0\] Now you can solve for x, and then use x = tan(theta)

OpenStudy (agent0smith):

Oops, 2nd to last line should be x+1 not x-1.

OpenStudy (anonymous):

i only got 60 degrees

OpenStudy (agent0smith):

x=sqrt3 and x=-1, and x=tan(theta). There's... looks like 4 solutions total for 0º≤θ<360º \[\tan \theta = \sqrt 3\] \[\tan \theta = -1\] Find them on the unit circle, http://jwilson.coe.uga.edu/emat6680fa06/crumley/unit/6_files/image010.jpg

OpenStudy (anonymous):

yup, there have 4 answers

OpenStudy (agent0smith):

Did you find all 4?

OpenStudy (agent0smith):

Just look for where tan = sqrt 3 and -1.

OpenStudy (anonymous):

nope, i just find 60degrees

OpenStudy (agent0smith):

Look harder :P

OpenStudy (agent0smith):

I can see four instances on here: http://jwilson.coe.uga.edu/emat6680fa06/crumley/unit/6_files/image010.jpg where tan is either equal to -1, or sqrt3.

OpenStudy (agent0smith):

Here's two:

OpenStudy (anonymous):

ummm...

OpenStudy (anonymous):

180+(-45)=135

OpenStudy (agent0smith):

Do you know how to use a unit circle? all the values you need are on it.

OpenStudy (anonymous):

no..

OpenStudy (agent0smith):

http://withfriendship.com/images/b/8803/Unit-circle-picture.gif It shows the angles, and the values of cosine and sine.

OpenStudy (agent0smith):

tan (x + 180°) = tan x you can also use this to find all of your solutions.

OpenStudy (agent0smith):

180+(-45)=135 with this 135 and 60 degrees, you can find them all.